Please show all steps.

(1 point) Solve the system 13 +2x4 +2x3 1 -21 24 -4x3 94 +163 12x4 +бх4 За1 —Зх2 41 +4x3 X2 2 III| ह
Answer #1

| 21 C2 | 31 -3 2 -I) +23 -423 +1623 +413 +24 = 13 - 24 = -21 +1214 = 94 +6x4 = 41 Augmented matun C 1 0 2 2 0 5 - 4 - 2 2 -- о R, → RitR3 , В, -, - - 2 o оо: o o о o o o о К. – К. Rч - Ru Г. | о o 2 - 2 Куә б o о т о о т o o - о о - R-2Rч Г. o К» Кplease rate